# A Comprehensive Survey on Knowledge Distillation

### v1

@@ IV-G Contrastive Distillation
where anchors ($F_{i}^{s}$) are typically features from the student’s embeddings, positives ($F_{i}^{t}$) are corresponding features from the teacher for the same input, and negatives ($F_{j}$) are unrelated features from other data points. These methods employ contrastive loss to align the student’s feature space with the teacher’s, enabling efficient knowledge transfer. Other studies [258, 259] innovate by introducing new contrastive loss functions that address challenges such as noisy negatives and capacity mismatches between teacher and student models, enhancing the robustness and effectiveness of the distillation process.
@@
Contrastive distillation has gained significant attention for its ability to enhance knowledge transfer through contrastive learning techniques. Tian et al.  [260] pioneered this approach by employing a contrastive loss to align the features of the teacher and student models, Expanding upon this foundation, WCoRD [261] introduced a method that incorporates both primal and dual forms of the Wasserstein distance. While the dual form ensures global knowledge transfer by maximizing mutual information, the primal form focuses on local feature alignment within mini-batches, providing a balanced approach to feature alignment and distillation.
In the field of semantic segmentation, [259] addresses the challenge of dense pixel-level representations by leveraging contrastive loss to align the teacher and student feature maps. Similarly, [262] focuses on the domain of single-image super-resolution. Their approach utilizes contrastive loss to distill statistical information from intermediate teacher features into a lightweight student network, enabling high-resolution image reconstruction with minimal computational resources.
The problem of distilling large language-image pretraining models, such as CLIP [263], is tackled by [264], which develops a novel framework incorporating image feature alignment and educational attention modules. This method effectively aligns multi-modal features through contrastive loss. To address challenges in sentence embedding, DistilCSE [258] proposes a method which compresses large contrastive sentence embedding models.
In the domain of medical imaging, CRCKD [265] combines class-guided contrastive distillation and categorical relation preserving techniques to enhance intra-class similarity and inter-class divergence. The use of class centroids and relational graphs ensures effective knowledge transfer, even in imbalanced datasets.
In the area of weakly supervised visual grounding, [266] uses pseudo labels generated by object detectors. By applying contrastive loss for region-phrase matching, their method eliminates the need for object detection during inference, simplifying the process. [267] tackles the problem of relation-based knowledge transfer with CRCD. By maximizing mutual information between anchor-teacher and anchor-student relation distributions, their framework aligns inter-sample relations through a relation contrastive loss.
@@ V Modalities
V-A 3D Input
Knowledge distillation techniques contribute significantly to enhancing various 3D-related tasks, including object detection, semantic segmentation, shape generation, and shape classification. This section explores the application of KD in 3D data, highlighting innovative approaches and categorizing contributions based on task domains. Figure 7 illustrates the key tasks, domains, and types of 3D data where KD techniques have been successfully applied. Table VI summarizes distillation methods for each task based on the source of distillation.
@@
LiDAR: Precise spatial and depth sensing are essential for autonomous systems, , yet challenges such as sparsity and occlusion often degrade performance in 3D object detection. To solve these, several KD frameworks have been proposed [268, 269, 270, 271, 272, 273, 274, 275, 276]. X-Ray Distillation [274] trains a teacher model on object-complete frames that are derived from aggregated LiDAR scans, to transfer knowledge to a student model that improves its ability to handle sparse data and occlusions. Similarly, RadarDistill [275] aligns radar and LiDAR features to handle sparsity and noise. Furthermore, RDD[271] reduces the mismatch between the teacher and student models by aligning feature representations and refining outputs.
In addition, PointDistiller  [272] focuses on local geometric structures in point clouds, using dynamic graph convolution and reweighted learning to focus on important points and voxels. itKD  [273] applies feature reduction and mutual refinement to transfer coarse and fine-grained geometric features. Additionally, SRKD  [276] bridges domain gaps caused by weather variations by aligning instance features through density and shape similarity and ensuring consistent predictions between the teacher and student models. LiDAR Distillation  [268] also addresses domain gaps caused by varying LiDAR beam densities by performing distillation from a high-beam LiDAR teacher to a low-beam student.
TABLE VI: Summary of 3D distillation methods based on their tasks and sources of distillation.
@@
Object Detection FD3D [277], X3KD [210], Gambashidze et al. [274], PointDistiller [272], itKD [273], CRKD [211], DistillBEV [209], RDD [271], RadarDistill [275], SRKD [276] FD3D [277], X3KD [210], Gambashidze et al. [274], itKD [273], CRKD [211] X3KD [210], Gambashidze et al. [274], PointDistiller [272], CRKD [211], DistillBEV [209], RDD [271]
Classification FAD [278], HSD [279] FAD [278], JGEKD [280], PointViG-Distil [281]
@@
Segmentation Liu et al. [206], Qiu et al. [282], CMDFusion [283], Jiang et al. [284], Smaller3d [285], Seal [286] PVD [287], Qiu et al. [282] Genova et al. [288], PSD [289], PVD [287], Qiu et al. [282], LGKD [290], Smaller3d [285], PartDistill [291]
Depth Estimation MVP-Net [292], LiRCDepth [293] LiRCDepth [293] MVP-Net [292], KD-MonoRec [294], LiRCDepth [293]
@@
Representation PPKT[295], Fu et al.[296], SLidR[297], PointCMT[298], RECON[299], MVNet[300], LiDAR2Map[301], C2P[302], HVDistill[303], Yao et al.[304], Text4Point[305], Diff3F[306] Yu et al.[307], LiDAR2Map[301], I2P-MAE[308]
Domain Adaptation Cardace et al. [309], Wu et al. [310], SEN [311] SEN [311]
Recognition DistilVPR [312], PointMCD [313]
@@
Completion SCPNet [314], RaPD [315], Lin et al. [316], VPNet [317], Huang et al. [318] ADNet [319] Hwang et al. [320], MonDi [321], Zhou et al. [322], Zhang et al. [323], HASSC [324], Huang et al. [318]
Registration Jiang et al. [325] DiReg [326]
Other 3D Tasks DTC123 [327], CSD [328], Van et al. [329] Van et al. [329] PCDNet [330]
@@
Moreover, another study [269] introduces two novel techniques: pivotal position logit KD, which focuses on key areas for enhanced distillation, and teacher-guided initialization, which transfers the teacher’s feature extraction capabilities to the student through weight inheritance. However, intrinsic challenges such as sparsity, randomness, and varying density limit the effectiveness of normal distillation algorithms. To overcome these limitations, PVD [270] leverages the advantages of both point-level and voxel-level knowledge to address these challenges in LiDAR object detection tasks.
@@
Camera-only (single or multi-camera): Achieving accurate 3D object detection using only cameras is challenging due to the absence of depth information. To address this, FD3D [277] uses selective masked generative distillation and query-based focal distillation to enhance object-specific learning in perspective and Bird’s-Eye View (BEV) spaces. Similarly, X3KD [210] integrates cross-task, cross-modal, and cross-stage distillation with techniques such as LiDAR feature alignment and adversarial training for dense supervision.
@@
Cross-modal: Transferring spatial and depth knowledge from LiDAR-based models to cost-effective camera-radar or camera-only alternatives enhances multi-sensor systems, improving detection performance while resolving sensor inconsistencies. CRKD  [211] transfers knowledge from LiDAR-camera teacher to a camera-radar student in the BEV space. Similarly, DistillBEV  [209] transfers 3D geometric knowledge from a LiDAR-based teacher to a multi-camera BEV student using region decomposition, adaptive scaling, spatial attention, and multi-scale distillation. In addition, CMKD  [207] transfers spatial knowledge from LiDAR-based teachers to monocular students, tackling depth estimation challenges using BEV features and teacher predictions. In addition UniDistill  [208] proposes a universal KD framework which supports multiple modality pairs (LiDAR-to-camera, camera-to-LiDAR, fusion-to-camera, and fusion-to-LiDAR). The model employed BEV as a shared representation to align teacher and student detectors across modalities.
V-A2 3D Shape Classification
@@
Point cloud classification is another critical task where the teacher-student framework can be applied. FAD [278] introduces a novel loss function that combines logit distillation and feature distillation. JGEKD [281] proposes a loss function based on joint graph entropy to address the challenges of non-independent and identically distributed 3D point cloud data in classification tasks. To improve the efficiency, Zhang et al. in [280] introduced an offline distillation framework that incorporates a negative-weight self-distillation approach. [279] employs a self-distillation framework for incomplete point cloud classification.
V-A3 3D Semantic Segmentation
@@
KD enhances the accuracy and robustness of point cloud segmentation by effectively transferring knowledge from a high-capacity teacher model to a lightweight student model [288, 206, 289, 287, 282, 290, 283, 284, 285, 124, 291, 286].
@@
In order to address challenges such as sparsity, randomness and varying density in segmentation, [287] proposes point-to-voxel KD which distills the probabilistic outputs of the teacher at both the point level (fine-grained details) and voxel level (coarse structural information). Qiu et al. [282] proposes a multi-to-single KD framework that fuses multi-scan information for hard classes and employs a multilevel distillation strategy, including feature, logit, and instance-aware similarity distillation.
@@
Limited labeled data is another major challenge in the semantic segmentation of large-scale point clouds. Despite the task of annotating pointwise labels for such datasets being expensive and time-consuming, existing methods often rely on fully supervised approaches that require dense annotations. PSD [289] addresses this issue by introducing perturbed branches and leveraging graph-based consistency in a self-distillation manner. Seal [286] is a self-supervised learning framework that distills knowledge from off-the-shelf vision foundation models for point cloud segmentation. PartDistill [291] proposes a method to distill knowledge from a vision-language model to a 3D segmentation network.
V-A4 3D Domain Adaptation
@@
Unsupervised Domain Adaptation (UDA) struggles to perform effectively when the target domain differs from the labeled source domain, due to factors such as noise, occlusions, or missing elements. [309] proposes a self-distillation UDA technique to generate discriminative representations for the target domain and utilizes GNN-based online pseudo-label refinement. [310] introduces a cross-modal feature fusion in UDA for semantic segmentation. [311] proposes a self-ensembling network for domain adaptation in 3D point clouds, which leverages a semi-supervised learning framework for effective knowledge transfer from a labeled source domain to an unlabeled target domain.
V-A5 3D Depth Estimation
@@
3D depth estimation is crucial for accurately understanding the geometry of a scene, enabling perception systems to reconstruct spatial relationships and object structures. MVP-Net [292] enhances reconstruction by using depth estimation in a multi-view, cross-modal distillation approach for better point cloud upsampling. KD-MonoRec [294] leverages monocular RGB images as input and employs KD along with point cloud optimization to improve depth estimation. LiRCDepth [293] utilizes RGB images and radar point clouds and proposes a lightweight depth estimation model via KD.
V-A6 3D Representation Learning
@@
Representation learning focuses on extracting meaningful representations from data. Collecting a large and high-quality annotated dataset in 3D is a costly and time-consuming process, and pre-training 3D models requires access to large-scale 3D datasets, which is not a trivial task. However, some works, such as DCGLR [296], have effectively utilized knowledge distillation (KD) with only 3D data to address this challenge. However, given that numerous large-scale datasets and powerful foundation models are available in the 2D domain, several methods [295, 297, 307, 298, 301, 308, 300, 304, 303, 306] have leveraged these capabilities to learn rich representations for 3D point clouds.
@@
[295] introduces contrastive learning to transfer 2D semantic knowledge to 3D networks. SLidR [297] introduces a self-supervised 2D-to-3D representation distillation framework that uses superpixel-driven contrastive loss to align image and LiDAR features, enabling effective pre-training of 3D perception models for autonomous driving. [299] proposes a framework that can be used in either a single-modal or cross-modal setting.
@@
3D representations can be learned with the help of text [305] or through sequences of point clouds, as demonstrated in [302], which develops a self-supervised method for learning 4D point cloud representations using KD.
V-A7 3D Recognition
@@
Visual place recognition is an important task in computer vision and robotics, aiming to identify previously visited locations based on visual input such as camera image and LiDAR point clouds. PointMCD [313] and DistilVPR [312] are two innovative cross-modal KD frameworks in this field, with PointMCD employing a multi-view framework, as discussed further in Section V-B.
V-A8 3D Completion
@@
Shape Completion: Shape completion aims to reconstruct a complete point cloud from occluded input point cloud. RaPD [315] is the first semi-supervised point cloud completion method that utilizes prior distillation. [316] uses loss distillation and [322] propose a hierarchical self-distillation point cloud completion method.
Scene Completion: LiDAR sensors often capture incomplete point clouds due to occlusions from objects or their surroundings, making large-scale 3D scene interpretation challenging. KD can enhance their performance as mentioned in [314, 324, 323, 317].
Depth Completion: Depth completion is the process of estimating a dense depth map from sparse or incomplete depth measurements, such as those obtained from LiDAR sensors. KD can aid in predicting missing depth information [320, 321, 319, 318].
V-A9 3D Registration
KD can enhance point cloud registration by transferring knowledge from a large, complex model to a smaller, efficient one while preserving performance [326, 325].
V-A10 Other 3D Tasks
KD is applied to various other 3D tasks with different types of distillation. This includes generation [327], stylization [328], sampling [330] and accordance detection, as in [329], which utilizes all types of KD.
V-B Multi-view Input

### v2

@@ 4.7 Contrastive Distillation
where anchors ($F_{i}^{s}$) are typically features from the student’s embeddings, positives ($F_{i}^{t}$) are corresponding features from the teacher for the same input, and negatives ($F_{j}$) are unrelated features from other data points. These methods employ contrastive loss to align the student’s feature space with the teacher’s, enabling efficient knowledge transfer. Other studies (Gao et al., 2021; Fan et al., 2023) innovate by introducing new contrastive loss functions that address challenges such as noisy negatives and capacity mismatches between teacher and student models, enhancing the robustness and effectiveness of the distillation process.
@@
Contrastive distillation has gained significant attention for its ability to enhance knowledge transfer through contrastive learning techniques. (Tian et al., 2019) pioneered this approach by employing a contrastive loss to align the features of the teacher and student models, Expanding upon this foundation, WCoRD (Chen et al., 2021b) introduced a method that incorporates both primal and dual forms of the Wasserstein distance. While the dual form ensures global knowledge transfer by maximizing mutual information, the primal form focuses on local feature alignment within mini-batches, providing a balanced approach to feature alignment and distillation.
In the field of semantic segmentation, Fan et al. (2023) addresses the challenge of dense pixel-level representations by leveraging contrastive loss to align the teacher and student feature maps. Similarly, Liu et al. (2024b) focuses on the domain of single-image super-resolution. Their approach utilizes contrastive loss to distill statistical information from intermediate teacher features into a lightweight student network, enabling high-resolution image reconstruction with minimal computational resources.
The problem of distilling large language-image pretraining models, such as CLIP (Radford et al., 2021), is tackled by Chen et al. (2024g), which develops a novel framework incorporating image feature alignment and educational attention modules. This method effectively aligns multi-modal features through contrastive loss. To address challenges in sentence embedding, DistilCSE (Gao et al., 2021) proposes a method which compresses large contrastive sentence embedding models.
In the domain of medical imaging, CRCKD (Xing et al., 2021) combines class-guided contrastive distillation and categorical relation preserving techniques to enhance intra-class similarity and inter-class divergence. The use of class centroids and relational graphs ensures effective knowledge transfer, even in imbalanced datasets.
In the area of weakly supervised visual grounding, Wang et al. (2021e) uses pseudo labels generated by object detectors. By applying contrastive loss for region-phrase matching, their method eliminates the need for object detection during inference, simplifying the process. Zhu et al. (2021b) tackles the problem of relation-based knowledge transfer with CRCD. By maximizing mutual information between anchor-teacher and anchor-student relation distributions, their framework aligns inter-sample relations through a relation contrastive loss. Most recently, CRD (Zhu et al., 2025) has leveraged contrastive learning to align class-wise semantics by preserving sample-wise similarity for intra-sample alignment, while capturing structural semantics for inter-sample contrast.
@@ 5 Modalities
5.1 3D Input
Knowledge distillation techniques contribute significantly to enhancing various 3D-related tasks, including object detection, semantic segmentation, shape generation, and shape classification. This section explores the application of KD in 3D data, highlighting innovative approaches and categorizing contributions based on task domains. Figure 7 illustrates the key tasks, domains, and types of 3D data where KD techniques have been successfully applied. Table 6 summarizes distillation methods for each task based on the source of distillation.
@@
LiDAR: Precise spatial and depth sensing are essential for autonomous systems, , yet challenges such as sparsity and occlusion often degrade performance in 3D object detection. To solve these, several KD frameworks have been proposed (Wei et al., 2022; Yang et al., 2022d; Zhang & Liu, 2023; Li et al., 2023g; Zhang et al., 2023d; Cho et al., 2023; Gambashidze et al., 2024; Bang et al., 2024; Huang et al., 2024e). X-Ray Distillation (Gambashidze et al., 2024) trains a teacher model on object-complete frames that are derived from aggregated LiDAR scans, to transfer knowledge to a student model that improves its ability to handle sparse data and occlusions. Similarly, RadarDistill (Bang et al., 2024) aligns radar and LiDAR features to handle sparsity and noise. Furthermore, RDD(Li et al., 2023g) reduces the mismatch between the teacher and student models by aligning feature representations and refining outputs.
In addition, PointDistiller  (Zhang et al., 2023d) focuses on local geometric structures in point clouds, using dynamic graph convolution and reweighted learning to focus on important points and voxels. itKD  (Cho et al., 2023) applies feature reduction and mutual refinement to transfer coarse and fine-grained geometric features. Additionally, SRKD  (Huang et al., 2024e) bridges domain gaps caused by weather variations by aligning instance features through density and shape similarity and ensuring consistent predictions between the teacher and student models. LiDAR Distillation  (Wei et al., 2022) also addresses domain gaps caused by varying LiDAR beam densities by performing distillation from a high-beam LiDAR teacher to a low-beam student.
Table 6: Summary of 3D distillation methods based on their tasks and sources of distillation.
@@
Object Detection FD3D (Zeng et al., 2023), X3KD (Klingner et al., 2023), Gambashidze et al. (Gambashidze et al., 2024), PointDistiller (Zhang et al., 2023d), itKD (Cho et al., 2023), CRKD (Zhao et al., 2024b), DistillBEV (Wang et al., 2023p), RDD (Li et al., 2023g), RadarDistill (Bang et al., 2024), SRKD (Huang et al., 2024e) FD3D (Zeng et al., 2023), X3KD (Klingner et al., 2023), Gambashidze et al. (Gambashidze et al., 2024), itKD (Cho et al., 2023), CRKD (Zhao et al., 2024b) X3KD (Klingner et al., 2023), Gambashidze et al. (Gambashidze et al., 2024), PointDistiller (Zhang et al., 2023d), CRKD (Zhao et al., 2024b), DistillBEV (Wang et al., 2023p), RDD (Li et al., 2023g)
Classification FAD (Lee & Wu, 2023), HSD (Zhou et al., 2024a) FAD (Lee & Wu, 2023), JGEKD (Zheng et al., 2024a), PointViG-Distil (Tian et al., 2023b)
@@
Segmentation Liu et al. (Liu et al., 2021g), Qiu et al. (Qiu et al., 2023), CMDFusion (Cen et al., 2023), Jiang et al. (Jiang et al., 2023a), Smaller3d (Adamyan & Harutyunyan, 2023), Seal (Liu et al., 2024h) PVD (Hou et al., 2022), Qiu et al. (Qiu et al., 2023) Genova et al. (Genova et al., 2021), PSD (Zhang et al., 2021f), PVD (Hou et al., 2022), Qiu et al. (Qiu et al., 2023), LGKD (Yang et al., 2023f), Smaller3d (Adamyan & Harutyunyan, 2023), PartDistill (Umam et al., 2024)
Depth Estimation MVP-Net (Chen et al., 2024b), LiRCDepth (Sun et al., 2025) LiRCDepth (Sun et al., 2025) MVP-Net (Chen et al., 2024b), KD-MonoRec (Xiao et al., 2024), LiRCDepth (Sun et al., 2025)
@@
Representation PPKT(Liu et al., 2021d), Fu et al.(Fu et al., 2022), SLidR(Sautier et al., 2022), PointCMT(Yan et al., 2022), RECON(Qi et al., 2023), MVNet(Yan et al., 2023), LiDAR2Map(Wang et al., 2023m), C2P(Zhang et al., 2023k), HVDistill(Zhang et al., 2024f), Yao et al.(Yao et al., 2024), Text4Point(Huang et al., 2024c), Diff3F(Dutt et al., 2024) Yu et al.(Yu et al., 2022b), LiDAR2Map(Wang et al., 2023m), I2P-MAE(Zhang et al., 2023g)
Domain Adaptation Cardace et al. (Cardace et al., 2023), Wu et al. (Wu et al., 2023e), SEN (Li et al., 2025a) SEN (Li et al., 2025a)
Recognition DistilVPR (Wang et al., 2024h), PointMCD (Zhang et al., 2023f)
@@
Completion SCPNet (Xia et al., 2023), RaPD (Fan et al., 2022), Lin et al. (Lin et al., 2024a), VPNet (Wang et al., 2025b), Huang et al. (Huang et al., 2024a) ADNet (Kim et al., 2024b) Hwang et al. (Hwang et al., 2021), MonDi (Liu et al., 2022b), Zhou et al. (Zhou et al., 2024b), Zhang et al. (Zhang et al., 2024e), HASSC (Wang et al., 2024i), Huang et al. (Huang et al., 2024a)
Registration Jiang et al. (Jiang et al., 2024a) DiReg (Löwens et al., 2024)
Other 3D Tasks DTC123 (Yi et al., 2024b), CSD (Decatur et al., 2024), Van et al. (Van Vo et al., 2024) Van et al. (Van Vo et al., 2024) PCDNet (Huang et al., 2022c)
@@
Moreover, another study (Yang et al., 2022d) introduces two novel techniques: pivotal position logit KD, which focuses on key areas for enhanced distillation, and teacher-guided initialization, which transfers the teacher’s feature extraction capabilities to the student through weight inheritance. However, intrinsic challenges such as sparsity, randomness, and varying density limit the effectiveness of normal distillation algorithms. To overcome these limitations, PVD (Zhang & Liu, 2023) leverages the advantages of both point-level and voxel-level knowledge to address these challenges in LiDAR object detection tasks.
@@
Camera-only (single or multi-camera): Achieving accurate 3D object detection using only cameras is challenging due to the absence of depth information. To address this, FD3D (Zeng et al., 2023) uses selective masked generative distillation and query-based focal distillation to enhance object-specific learning in perspective and Bird’s-Eye View (BEV) spaces. Similarly, X3KD (Klingner et al., 2023) integrates cross-task, cross-modal, and cross-stage distillation with techniques such as LiDAR feature alignment and adversarial training for dense supervision.
@@
Cross-modal: Transferring spatial and depth knowledge from LiDAR-based models to cost-effective camera-radar or camera-only alternatives enhances multi-sensor systems, improving detection performance while resolving sensor inconsistencies. CRKD  (Zhao et al., 2024b) transfers knowledge from LiDAR-camera teacher to a camera-radar student in the BEV space. Similarly, DistillBEV  (Wang et al., 2023p) transfers 3D geometric knowledge from a LiDAR-based teacher to a multi-camera BEV student using region decomposition, adaptive scaling, spatial attention, and multi-scale distillation. In addition, CMKD  (Hong et al., 2022b) transfers spatial knowledge from LiDAR-based teachers to monocular students, tackling depth estimation challenges using BEV features and teacher predictions. In addition UniDistill  (Zhou et al., 2023b) proposes a universal KD framework which supports multiple modality pairs (LiDAR-to-camera, camera-to-LiDAR, fusion-to-camera, and fusion-to-LiDAR). The model employed BEV as a shared representation to align teacher and student detectors across modalities.
5.1.2 3D Shape Classification
@@
Point cloud classification is another critical task where the teacher-student framework can be applied. FAD (Lee & Wu, 2023) introduces a novel loss function that combines logit distillation and feature distillation. JGEKD (Tian et al., 2023b) proposes a loss function based on joint graph entropy to address the challenges of non-independent and identically distributed 3D point cloud data in classification tasks. To improve the efficiency Zheng et al. (2024a) introduced an offline distillation framework that incorporates a negative-weight self-distillation approach. Zhou et al. (2024a) employs a self-distillation framework for incomplete point cloud classification.
5.1.3 3D Semantic Segmentation
@@
KD enhances the accuracy and robustness of point cloud segmentation by effectively transferring knowledge from a high-capacity teacher model to a lightweight student model (Genova et al., 2021; Liu et al., 2021g; Zhang et al., 2021f; Hou et al., 2022; Qiu et al., 2023; Yang et al., 2023f; Cen et al., 2023; Jiang et al., 2023a; Adamyan & Harutyunyan, 2023; Karine et al., 2024; Umam et al., 2024; Liu et al., 2024h).
@@
In order to address challenges such as sparsity, randomness and varying density in segmentation, Hou et al. (2022) proposes point-to-voxel KD which distills the probabilistic outputs of the teacher at both the point level (fine-grained details) and voxel level (coarse structural information). Qiu et al. (2023) proposes a multi-to-single KD framework that fuses multi-scan information for hard classes and employs a multilevel distillation strategy, including feature, logit, and instance-aware similarity distillation.
@@
Limited labeled data is another major challenge in the semantic segmentation of large-scale point clouds. Despite the task of annotating point-wise labels for such datasets being expensive and time-consuming, existing methods often rely on fully supervised approaches that require dense annotations. PSD (Zhang et al., 2021f) addresses this issue by introducing perturbed branches and leveraging graph-based consistency in a self-distillation manner. Seal (Liu et al., 2024h) is a self-supervised learning framework that distills knowledge from off-the-shelf vision foundation models for point cloud segmentation. PartDistill (Umam et al., 2024) proposes a method to distill knowledge from a vision-language model to a 3D segmentation network.
5.1.4 3D Domain Adaptation
@@
Unsupervised Domain Adaptation (UDA) struggles to perform effectively when the target domain differs from the labeled source domain, due to factors such as noise, occlusions, or missing elements. Cardace et al. (2023) proposes a self-distillation UDA technique to generate discriminative representations for the target domain and utilizes GNN-based online pseudo-label refinement. Wu et al. (2023e) introduces a cross-modal feature fusion in UDA for semantic segmentation. Li et al. (2025a) proposes a self-ensembling network for domain adaptation in 3D point clouds, which leverages a semi-supervised learning framework for effective knowledge transfer from a labeled source domain to an unlabeled target domain.
5.1.5 3D Depth Estimation
@@
3D depth estimation is crucial for accurately understanding the geometry of a scene, enabling perception systems to reconstruct spatial relationships and object structures. MVP-Net (Chen et al., 2024b) enhances reconstruction by using depth estimation in a multi-view, cross-modal distillation approach for better point cloud upsampling. KD-MonoRec (Xiao et al., 2024) leverages monocular RGB images as input and employs KD along with point cloud optimization to improve depth estimation. LiRCDepth (Sun et al., 2025) utilizes RGB images and radar point clouds and proposes a lightweight depth estimation model via KD.
5.1.6 3D Representation Learning
@@
Representation learning focuses on extracting meaningful representations from data. Collecting a large and high-quality annotated dataset in 3D is a costly and time-consuming process, and pre-training 3D models requires access to large-scale 3D datasets, which is not a trivial task. However, some works, such as DCGLR (Fu et al., 2022), have effectively utilized knowledge distillation (KD) with only 3D data to address this challenge. However, given that numerous large-scale datasets and powerful foundation models are available in the 2D domain, several methods (Liu et al., 2021d; Sautier et al., 2022; Yu et al., 2022b; Yan et al., 2022; Wang et al., 2023m; Zhang et al., 2023g; Yan et al., 2023; Yao et al., 2024; Zhang et al., 2024f; Dutt et al., 2024) have leveraged these capabilities to learn rich representations for 3D point clouds.
@@
Liu et al. (2021d) introduces contrastive learning to transfer 2D semantic knowledge to 3D networks. SLidR (Sautier et al., 2022) introduces a self-supervised 2D-to-3D representation distillation framework that uses superpixel-driven contrastive loss to align image and LiDAR features, enabling effective pre-training of 3D perception models for autonomous driving. Qi et al. (2023) proposes a framework that can be used in either a single-modal or cross-modal setting.
@@
3D representations can be learned with the help of text (Huang et al., 2024c) or through sequences of point clouds, as demonstrated in Zhang et al. (2023k), which develops a self-supervised method for learning 4D point cloud representations using KD.
5.1.7 3D Recognition
@@
Visual place recognition is an important task in computer vision and robotics, aiming to identify previously visited locations based on visual input such as camera image and LiDAR point clouds. PointMCD (Zhang et al., 2023f) and DistilVPR (Wang et al., 2024h) are two innovative cross-modal KD frameworks in this field, with PointMCD employing a multi-view framework, as discussed further in Section 5.2.
5.1.8 3D Completion
@@
Shape Completion: Shape completion aims to reconstruct a complete point cloud from occluded input point cloud. RaPD (Fan et al., 2022) is the first semi-supervised point cloud completion method that utilizes prior distillation. Lin et al. (2024a) uses loss distillation and Zhou et al. (2024b) propose a hierarchical self-distillation point cloud completion method.
Scene Completion: LiDAR sensors often capture incomplete point clouds due to occlusions from objects or their surroundings, making large-scale 3D scene interpretation challenging. KD can enhance their performance as mentioned in Xia et al. (2023); Wang et al. (2024i); Zhang et al. (2024e); Wang et al. (2025b).
Depth Completion: Depth completion is the process of estimating a dense depth map from sparse or incomplete depth measurements, such as those obtained from LiDAR sensors. KD can aid in predicting missing depth information (Hwang et al., 2021; Liu et al., 2022b; Kim et al., 2024b; Huang et al., 2024a).
5.1.9 3D Registration
KD can enhance point cloud registration by transferring knowledge from a large, complex model to a smaller, efficient one while preserving performance (Löwens et al., 2024; Jiang et al., 2024a).
5.1.10 Other 3D Tasks
KD is applied to various other 3D tasks with different types of distillation. This includes generation (Yi et al., 2024b), stylization (Decatur et al., 2024), sampling (Huang et al., 2022c) and accordance detection, as in (Van Vo et al., 2024), which utilizes all types of KD.
5.2 Multi-view Input